# Neural Posterior Estimation for Cataloging Astronomical Images with Spatially Varying Backgrounds and Point Spread Functions

### v1

@@ Acknowledgments
Our software is available in a public GitHub repository: https://github.com/prob-ml/bliss. The specific code for reproducing the figures and tables in this article is within the case_studies/psf_variation directory.
@@ Appendix A Performance stratified by Signal-to-Noise Ratio
In this section, we repeat our evaluation from Section 7.1 but report performance stratified by the signal-to-noise ratios (SNRs) of objects rather than the magnitudes of the objects. The SNR of an object depends on the flux within the object’s footprint, noise from the background, and Poisson noise of the object. We determine the footprint by masking out all pixels whose fluxes did not appreciably change with the addition of a new object. Then, following Jones (2024), we compute the SNR as follows:
@@
Figure 9: Performance for held-out data with a varied backgrounds and PSFs. This figure displays the data from Figure 4 with performance stratified by signal-to-noise (SNR) rather than object magnitude.
Figure 9 is analogous to Figure 4 in the main text, with performance stratified by signal-to-noise (SNR) rather than object magnitude. Precision for light source detection is omitted from Figure 9 because it is unclear how to compute the SNR of unmatched detections.
Like Figure 4, Figure 9 shows 1) nearly monotonic performance of the PSF-unaware and PSF-aware networks, 2) the single-field network often performing worse than the PSF-unaware and PSF-aware networks without ever performing better, and 3) the PSF-aware network sometimes outperforming the PSF-unaware network without ever performing worse.

### v2

@@ Acknowledgments
Our software is available in a public GitHub repository at https://github.com/prob-ml/bliss and a static version is posted on Zenodo (Patel et al., 2025). The specific code for reproducing the figures and tables in this article is within the case_studies/psf_variation directory.
@@ Appendix A Performance stratified by Signal-to-Noise Ratio
In this appendix, we repeat our evaluation from Section 6.1 but report performance stratified by the signal-to-noise ratios (SNRs) of objects rather than the magnitudes of the objects. The SNR of an object depends on the flux within the object’s footprint, noise from the background, and Poisson noise of the object. We determine the footprint by masking out all pixels whose fluxes did not appreciably change with the addition of a new object. Then, following Jones (2024), we compute the SNR as follows:
@@
Figure A.1: Performance for held-out data with varied backgrounds and PSFs. This figure displays the data from Figure 4 with performance stratified by signal-to-noise (SNR) rather than object magnitude.
Figure A.1 is analogous to Figure 4 in the main text, with performance stratified by signal-to-noise (SNR) rather than object magnitude. Precision for light source detection is omitted from Figure A.1 because it is unclear how to compute the SNR of unmatched detections.
Like Figure 4, Figure A.1 shows 1) nearly monotonic performance of the PSF-unaware and PSF-aware networks, 2) the single-field network often performing worse than the PSF-unaware and PSF-aware networks without ever performing better, and 3) the PSF-aware network sometimes outperforming the PSF-unaware network without ever performing worse.
Appendix B Image Reconstruction
In Figure B.1, we show samples of astronomical images from a held-out semi-synthetic dataset and the catalogs of them produced by our inference network.
Figure B.1: Three examples of cataloging with the “PSF-aware” inference network, one per row. The left column shows input images, which are held-out and semi-synthetic, along with the true positions and types (star vs galaxy) of the imaged astronomical objects. The center column show our predictions of the positions and types of astronomical objects in these images (that is, mode of variational distribution), as well as model reconstructions of the images. The right column shows the residual for each pixel.